OpenStudy (anonymous):

One leg of a right triangle is 6 inches longer than the other leg. The hypotenuse of the triangle is 25 inches. What is the length of each leg to the nearest inch? Separate answers using commas.

OpenStudy (anonymous):

@shamil98

OpenStudy (anonymous):

do you know the pythagorean theorem? Asquared + Bsquared=C squared, C being the hypotenuse.

OpenStudy (anonymous):

\[x ^{2} + (x+6)^{2} = 25^{2}\]

OpenStudy (austinl):



OpenStudy (austinl):

Pythagorean Theorem, \(a^2+b^2=c^2\) Where a and b are the legs, and c is the hypotenuse. \(\therefore x^2+(x+6)^2=25^2\)

OpenStudy (austinl):

When you do the math for the left side, you have to FOIL out that sum of x+6. \((x+6)^2=(x+6)(x+6)=x^2+12x+36\) So inputting, \(x^2+x^2+12x+36=625\) Combine like terms, \(2x^2+12x+36=625\) Then you can move the 625 over so that you have a nice quadratic equation. \(2x^2+12x-589=0\) Now, you can use the quadratic equation, \(\displaystyle x=\frac{-b\pm\sqrt{b^2-4ac}}{2a},~\text{when}~ax^2+bx-c=0\)

OpenStudy (anonymous):

so then what would the lengths be?

OpenStudy (austinl):

You should be able to do that yourself! It is spelled out for you above :)

OpenStudy (anonymous):

sorry, I didnt see it, my computer wasn't showing it, but I refreshed the page and now it's all good! thanks!

OpenStudy (austinl):

No problemo!

OpenStudy (anonymous):

so when I put it in the quadratic formula, this is the work I did: \[2x^2 + 12x - 589 = 0\] \[x = \frac{-(12)^2 \pm \sqrt{(12)^2 - 4(2)(-589)}}{ 2(2) }\] \[x = \frac{-144 \pm \sqrt{144 - 4,712} }{ 4 }\] then what would I do from there @austinL

OpenStudy (austinl):

You made an error, it is -b not -b^2

OpenStudy (anonymous):

where at?

OpenStudy (austinl):

\(\displaystyle x=\frac{\color{red}{-b}\pm\sqrt{b^2-4ac}}{2a},~\text{when}~ax^2+bx-c=0\)

OpenStudy (anonymous):

oh oh okay! so then it's \[x = \frac{ -12 \pm \sqrt{144 - 4,712} }{ 4 }\]

OpenStudy (austinl):

\(\displaystyle x = \frac{ -12 \pm \sqrt{144 - (-4,712)} }{ 4 }\)

OpenStudy (anonymous):

\[x = \frac{ -12 \pm \sqrt{4,856} }{ 4 }\]

OpenStudy (austinl):

Correctomundo. You will get two answers, one is impossible because it would make a leg longer than the hypotenuse. So as a result you should get, \(x=\sqrt{\dfrac{607}{2}}-3\approx 14.421\) Then the other leg is x+6

OpenStudy (anonymous):

okay, so how would i find the last leg?

OpenStudy (austinl):

\(x=\sqrt{\dfrac{607}{2}}-3\) \(x+6=(\sqrt{\dfrac{607}{2}}-3)+6=\sqrt{\dfrac{607}{2}}+3\)

OpenStudy (anonymous):

well you have x and x + 6 would just be 14.421 + 6.

OpenStudy (anonymous):

so all the legs are 25, 14.421, and 20.421?

OpenStudy (anonymous):

all inches of course.

OpenStudy (austinl):

No, legs are the decimals. Hypotenuse is 25.

OpenStudy (anonymous):

oh right, okay! you guys helped so much!

OpenStudy (austinl):

My pleasure! and good job!

OpenStudy (anonymous):

thank you:)